# Neutron-nucleus dynamics simulations for quantum computers

### 22 Feb 2024 v1

@@ Abstract
With a view toward addressing the explosive growth in the computational demands of nuclear structure and reactions modeling, we develop a novel quantum algorithm for neutron-nucleus simulations with general potentials, which provides acceptable bound-state energies even in the presence of noise, through the noise-resilient training method. In particular, the algorithm can now solve for any band-diagonal to full Hamiltonian matrices, as needed to accommodate a general central potential. This includes exponential Gaussian-like potentials and ab initio inter-cluster potentials (optical potentials). The approach can also accommodate the complete form of the chiral effective-field-theory nucleon-nucleon potentials used in ab initio nuclear calculations. We make this potential available for three different qubit encodings, including the one-hot, binary, and Gray encodings, and we provide a comprehensive analysis of the number of Pauli terms and commuting sets involved. We find that the Gray encoding allows for an efficient scaling of the model-space size $N$ (or number of basis states used) and is more resource efficient not only for tridiagonal Hamiltonians, as suggested earlier, but also for band-diagonal Hamiltonians having bandwidth up to $N$. We introduce a new commutativity scheme called distance-grouped commutativity (DGC) and compare its performance with the well-known qubit-commutativity (QC) scheme. We lay out the explicit grouping of Pauli strings and the diagonalizing unitary under the DGC scheme, and we find that it outperforms the QC scheme, at the cost of a more complex diagonalizing unitary. Lastly, we provide first solutions of the neutron-alpha dynamics from quantum simulations suitable for noisy intermediate-scale quantum processors, using an optical potential rooted in first principles, as well as a study of the bound-state physics in neutron-Carbon systems, along with a comparison of the efficacy of the one-hot and Gray encodings.
@@ I Introduction
With a view toward addressing such challenges in the long term by harnessing the advantages of quantum computing – as demonstrated for various low- and high-energy nuclear physics problems (e.g., see [13, 14, 15, 16, 17, 18, 19, 20, 21]) – in this paper, we start with the simplest case of two clusters, one of which is a neutron. We provide, for the first time, solutions of the neutron-nucleus dynamics from quantum simulations suitable for the far-term error-corrected regime as well as for the noisy intermediate-scale quantum (NISQ) processors coupled with the noise-resilient (NR) training method [22]: this is illustrated for the bound-state physics of the neutron-alpha (n-${}^{4}$e) optical potential rooted in first principles [23], as well as for the lowest $\frac{1}{2}^{+}$ energy in Carbon isotopes calculated through the n+${}^{10}$, n+${}^{12}$, and n+${}^{14}$ dynamics. We note that, in distinction to terminology often used in quantum computing, “dynamics simulation” or “simulation of (nuclear) dynamics” refer here to the problem of modeling the nuclear multi-cluster system using a specific inter-cluster potential.
@@
The present method utilizes the Variational Quantum Eigensolver (VQE) [24, 25, 26] and is based on the pioneering nuclear simulations of the deuteron on quantum computers, where the potential is given only by a single matrix element [27]. In our study, we design a novel quantum algorithm for a two-cluster system and a very general potential, which allows for versatile applications, including widely used exponential potentials in reaction calculations (see, e.g., Ref. [28]) and most importantly, optical potentials derived ab initio.
@@
Our algorithm makes these potentials available for three mappings to qubits: one-hot, binary, and Gray encodings (see, e.g., [29]). We explore advantages of each of these encodings and show that the Gray encoding (introduced to nuclear calculations in Ref. [30]) allows for an efficient scaling of the model-space size $N$ (or number of the basis states used) and is more resource efficient not only for tridiagonal Hamiltonians ($K=1$), as suggested in Ref. [29], but also for band-diagonal Hamiltonians for $K<N/2$, where $12K+$ 1 is the bandwidth of the Hamiltonian. Interestingly, we show that for bandwidths larger than $N$, more off-diagonals can be added, if needed for an increased accuracy, without increasing the complexity of the problem. Another outcome of this study relates to the efficacy of measurements, which is of key importance to obtaining acceptable outcomes on quantum devices. In particular, we introduce a new commutativity scheme called distance-grouped commutativity (DGC), which is especially useful for band-diagonal matrices. We compare its performance with the well-known qubit-commutativity (QC) scheme. We lay out the explicit grouping of Pauli strings and the diagonalizing unitary under the DGC scheme. We show that the DGC scheme outperforms the QC scheme, at the cost of a more complex diagonalizing unitary.
@@
We note that, in this study, the quantum simulations are reported for the lowest bound states of two clusters, for which a manageable number of qubits can be currently used (three or four qubits). Ultimately, the algorithm presented here can underpin multi-cluster dynamics simulations at low energies, such as those relevant to astrophysical studies, and can be utilized for weakly-bound states (e.g., for n+${}^{16}$ and n+${}^{18}$) and even for isolated low-lying resonances that require solutions in much larger model spaces (larger $N$) [^1] Applications of the present quantum algorithm are shown here for the use of harmonic oscillator single-particle basis states, the same square-integrable basis utilized in Refs. [31, 32, 23, 10], while the asymptotics are recovered in an R-matrix technique [12]. Alternatively, one can use the quantum algorithm for a square-nonintegrable basis, such as the eigenstates of the Woods-Saxon potential as utilized, e.g., in Ref. [33].. Knowledge about the bound-state physics is key, e.g., to the description of deuteron break-up reactions, such as (d,p) and (d,n) reactions, for which standard distorted-wave Born approximation methods rely on the physics of the bound states of the proton-nucleus and neutron-nucleus systems. As another important implication, exploring trade-offs of band-diagonal and full Hamiltonian matrices for different encodings is critical for the simplest two-cluster system of two nucleons; namely, this allows quantum simulations of nuclear structure to handle the complete form of the chiral nucleon-nucleon (NN) potentials (e.g., see [3, 4, 5]), which are in turn key to ab initio large-scale simulations of light, medium-mass, and even selected heavy nuclei.
We provide solutions for various n$+$C systems using an exponential potential, and for n$+\alpha$ based on the ab initio optical potential derived in Ref. [23]. We find that the quantum simulations are successful in finding the lowest $\frac{1}{2}^{+}$ bound-state energies. We develop a warm-start algorithm, inspired by perturbation theory, that is particularly useful for band-diagonal Hamiltonians. In this method, we first simulate the system for a simpler system, and use the endpoint of the simulation as the start for the full-scale simulation. We find that this method allows for a quicker convergence to the true energy value.
@@
Our paper is intended to serve interdisciplinary research at the intersection of nuclear physics and quantum information science, and in some cases includes details that may be well known in one of the fields but are pedagogical for researchers of the other field: our aim here is to provide a complete framework for the problem at hand. Our paper is organized as follows. In Sec. II, we introduce the nuclear problem of solving the neutron-nucleus dynamics and its Hamiltonian. In Sec. III, we discuss different encoding methods of mapping the given Hamiltonian to a form that can be simulated on a quantum computer. The different encoding methods we discuss include the one-hot encoding, binary encoding, and the Gray encoding. In Sec. IV, we briefly explain the variational principle, for completeness of presentation. Since the variational principle depends on the choice of a trial state, called ansatz, we delineate the different ansatz choices for the different encodings. In Sec. V, we analyze various advantages between the different encodings considered, including the number of Pauli terms and the number of commuting sets, for a most general local potential and its band-diagonal approximation. As part of our trade-off analysis for the different encodings of Hamiltonians, we introduce the new DGC measurement scheme to group Pauli strings into commuting operator sets. For this scheme we provide the explicit diagonalizing unitary and an analysis of the number of commuting sets. In Sec. VI, we provide quantum simulations for different encodings and nuclear systems. We discuss the results and challenges for weakly bound states.
@@ II Problem description
[…]
@@
where $n_{r},n_{r}^{\prime}$ are the relative harmonic oscillator radial node number. The matrix elements of $T$ for every $\ell$ are given by
$\langle n_{r}^{\prime}\ell|T|n_{r}\ell\rangle\coloneqq\frac{\hbar\omega}{2}\Bigg{[}\bigg{(}2n_{r}+\frac{3}{2}\bigg{)}\delta_{n_{r}^{\prime},n_{r}}\\ \qquad-\sqrt{\bigg{(}n_{r}-\frac{\ell}{2}\bigg{)}\bigg{(}n_{r}+\frac{\ell+1}{2}\bigg{)}}\delta_{n_{r}^{\prime},n_{r}-1}\\ \qquad-\sqrt{\bigg{(}n_{r}-\frac{\ell-2}{2}\bigg{)}\bigg{(}n_{r}+\frac{\ell+3}{2}\bigg{)}}\delta_{n_{r}^{\prime},n_{r}+1}\Bigg{]},$ (9)
@@
$\langle n_{r}^{\prime}\ell|r^{2}|n_{r}\ell\rangle\coloneqq b_{s}^{2}\Bigg{[}\bigg{(}2n_{r}+\frac{3}{2}\bigg{)}\delta_{n_{r}^{\prime},n_{r}}\\ \qquad+\sqrt{\bigg{(}n_{r}-\frac{\ell}{2}\bigg{)}\bigg{(}n_{r}+\frac{\ell+1}{2}\bigg{)}}\delta_{n_{r}^{\prime},n_{r}-1}\\ \qquad+\sqrt{\bigg{(}n_{r}-\frac{\ell-2}{2}\bigg{)}\bigg{(}n_{r}+\frac{\ell+3}{2}\bigg{)}}\delta_{n_{r}^{\prime},n_{r}+1}\Bigg{]},$ (12)
@@
Earlier quantum computational simulations of nuclear systems [27, 30] have used a contact potential for $\ell=0$, or $\left\langle n_{r}^{\prime}\right|V\left|n_{r}\right\rangle=V_{0}\delta_{n_{r},n_{r}^{\prime}}\delta_{n_{r},0}$ (Fig. 1a). In this work, the Hamiltonian is generalized for any band-diagonal to full matrix and can now accommodate a general central potential, such as exponential Gaussian-like potentials using Eq. (13) (e.g., [28]), along with ab initio inter-cluster potentials (as those derived in Ref. [23]) and the central part of chiral NN potentials using Eq. (7). In general, chiral NN potentials, such as in Refs. [3, 4, 5], used in ab initio nuclear calculations require, in addition, spin and isospin degrees of freedom $\alpha=n_{r}(\ell\frac{1}{2})jmt_{z}$, as described above. Including the additional spin-isospin quantum numbers leads to a larger set of basis states, but nonetheless, in the present framework this is straightforward by generalizing Eq. (8) to
@@
where $\langle\alpha^{\prime}|V|\alpha\rangle$ are known matrix elements for any chiral NN potential and the enumerating index $\alpha$ replaces $n_{r}$ in the mappings to qubits discussed next. Importantly, physically relevant potentials with a comparatively soft core or widely used potentials renormalized using, e.g., the Similarity Renormalization Group (SRG) technique [38], are band-diagonal as a result of the decoupling of low- and high-momentum configurations. Hence, the advantages found in this study as a function of the bandwidth $12K+$ 1 directly generalize to the complete form of the chiral NN potentials and their routinely used band-diagonal structure.
@@ III Mapping onto a quantum computer
Mappings from operators in the Fock space to Pauli strings are called encodings. Multiple encodings can be found in existing literature, for example, the one-hot encoding (OH), binary encoding, and the Gray encoding (see, e.g., [29, 30]).
In this work, we focus on these three encodings and analyze various properties and trade-offs. To provide an illustrative example, for each encoding above, we show the explicit Pauli terms for the $N=4,K=2$ Hamiltonian with the exponential potential in Eq. (13) with parameters
$V_{0}=-2.79\,{\rm MeV},\,c=0.05,\,\hbar\omega=15.95\,{\rm MeV},$ (17)
for the case of n+${}^{16}$ with reduced mass $\mu=\frac{16}{17}m_{\rm N}$ (here we use $m_{\rm N}=938.272029$ MeV for both protons and neutrons). The Hamiltonians for the different encodings can be found in their respective sections below.
@@ III.0.1 One-hot encoding
Next, we define the number operators and ladder operators. The number operator $|m\rangle\!\langle m|$ maps $|m\rangle$ to itself and maps all other basis states to 00:
@@
In this encoded basis, this action involves flipping the (m±i)plus-or-$(m\pm i)$ qubit to 00 and the $m$ qubit to 1:
@@
The upper limit of the $k$ summation is $\max(K,1)$ due to the fact that even if $K=0$, the first off-diagonal term in the Hamiltonian is non-zero as a result of the kinetic energy having two off-diagonal terms. Note that, to preserve the size of the matrix to be $N$, the summation over $k$ terminates if $m+k>N-1$. Thus, in the encoded one-hot basis, if $K>1$, the resulting matrix is $(2K+1)$-diagonal.

### 15 Mar 2026 v3

# thanks: Corresponding author

@@ Abstract
With a view toward addressing the explosive growth in the computational demands of nuclear structure and reactions modeling, we develop a novel quantum algorithm for neutron-nucleus simulations with general potentials, which provides acceptable bound-state energies even in the presence of noise, through the noise-resilient training method. In particular, the algorithm can now solve for any band-diagonal to full Hamiltonian matrices, as needed to accommodate a general central potential. While we illustrate the approach for exponential Gaussian-like potentials and ab initio inter-cluster potentials (optical potentials), it can also accommodate the complete form of the chiral effective-field-theory nucleon-nucleon potentials used in ab initio nuclear calculations. In this study, we provide a comprehensive analysis for the efficacy of this approach for three different qubit encodings, including the one-hot, binary, and Gray encodings, in terms of the number of Pauli strings and commuting sets involved. We also discuss the advantages of the algorithm for Hamiltonians of various band-diagonal widths, especially critical for potentials of perturbative nature, leading to a drastically reduced runtime of quantum simulations. We prove that the Gray encoding allows for an efficient scaling of the model-space size $N$ (or number of basis states used) and is more resource efficient for band-diagonal Hamiltonians having bandwidth up to $N$. We introduce a new commutativity scheme called distance-grouped commutativity (DGC) and compare its performance with the well-known qubit-commutativity (QC) scheme. We lay out the explicit grouping of Pauli strings and the diagonalizing unitary under the DGC scheme, and we prove that it outperforms the QC scheme, at the cost of a more complex diagonalizing unitary. Lastly, we provide first solutions of the neutron-alpha dynamics from quantum simulations suitable for noisy intermediate-scale quantum processors, using an optical potential rooted in first principles, as well as a study of the bound-state physics in neutron-Carbon systems, along with a comparison of the efficacy of the one-hot and Gray encodings.
@@ I Introduction
With a view toward addressing such challenges in the long term by harnessing the advantages of quantum computing – as demonstrated for various low- and high-energy nuclear physics problems (e.g., see [18, 13, 52, 48, 31, 34, 29, 51, 56, 14, 37, 6]) – in this paper, we start with the simplest case of two clusters, one of which is a neutron. We provide, for the first time, solutions of the neutron-nucleus dynamics from quantum simulations suitable for the far-term error-corrected regime as well as for the noisy intermediate-scale quantum (NISQ) processors coupled with the noise-resilient (NR) training method [44]: this is illustrated for the bound-state physics of the neutron-alpha (n-4He) optical potential rooted in first principles [11], as well as for the lowest $\frac{1}{2}^{+}$ energy in Carbon isotopes calculated through the n+10C, n+12C, and n+14C dynamics. We note that, in distinction to terminology often used in quantum computing, “dynamics simulation” or “simulation of (nuclear) dynamics” refer here to the problem of modeling the nuclear multi-cluster system using a specific inter-cluster potential.
@@
The present method utilizes the Variational Quantum Eigensolver (VQE) [38, 12, 5] and is based on the pioneering nuclear simulations of the deuteron on quantum computers, where the potential is given only by a single matrix element [18]. In our study, we design a novel quantum algorithm for a two-cluster system and a very general potential, which allows for versatile applications, including widely used exponential potentials in reaction calculations (see, e.g., Ref. [15]) and most importantly, optical potentials derived ab initio. We note here that the choice of VQE can be replaced with any other technique, for example, tensor network warm starting, etc.
@@
In this paper, we provide a generalized, extensible, and strong mathematical formulation of three mappings to qubits: one-hot, binary, and Gray encodings (see, e.g., [43]), explore their relative advantages, and illustrate these for simulations with the above-mentioned potentials. Going beyond the scope of earlier work that explored specific properties of mappings based on simulations only (e.g., see [46]), we provide mathematical proofs of scaling for these three encodings. Furthermore, the techniques that we use to prove these results are applicable for general encodings. Based on this, we show that the Gray encoding (introduced to nuclear calculations in Ref. [16]) allows for an efficient scaling of the model-space size $N$ (or number of the basis states used) and is more resource efficient not only for tridiagonal Hamiltonians ($K=1$), as suggested in Ref. [43], but also for band-diagonal Hamiltonians for $K<N/2$, where $2K+1$ is the bandwidth of the Hamiltonian. Interestingly, we show that for bandwidths larger than $N$, more off-diagonals can be added, if needed for an increased accuracy, without increasing the complexity of the problem. Another outcome of this study relates to the efficacy of measurements, which is of key importance to obtaining acceptable outcomes on quantum devices. In particular, we introduce a new commutativity scheme called distance-grouped commutativity (DGC), which is especially useful for band-diagonal matrices. We compare its performance with the well-known qubit-commutativity (QC) scheme. We lay out the explicit grouping of Pauli strings and the diagonalizing unitary under the DGC scheme. We show that the DGC scheme outperforms the QC scheme, at the cost of a more complex diagonalizing unitary. We note here that the diagonalizing unitary turns out to be the $\operatorname{GHZ}$ preparation unitary (See Lemma A.17 for the mathematical definition), and these have been used for diagonalization in quantum simulations [49, 25].
@@
We note that, in this study, the quantum simulations are reported for the lowest bound states of two clusters, for which a manageable number of qubits can be currently used (three or four qubits). Ultimately, the algorithm presented here can underpin multi-cluster dynamics simulations at low energies, such as those relevant to astrophysical studies, and can be utilized for weakly-bound states (e.g., for n+16C and n+18C) and even for isolated low-lying resonances that require solutions in much larger model spaces (larger NN) [^1] Applications of the present quantum algorithm are shown here for the use of harmonic oscillator single-particle basis states, the same square-integrable basis utilized in Refs. [39, 17, 11, 42], while the asymptotics are recovered in an R-matrix technique [50]. Alternatively, one can use the quantum algorithm for a square-nonintegrable basis, such as the eigenstates of the Woods-Saxon potential as utilized, e.g., in Ref. [36].. Knowledge about the bound-state physics is key, e.g., to the description of deuteron break-up reactions, such as (d,p) and (d,n) reactions, for which standard distorted-wave Born approximation methods rely on the physics of the bound states of the proton-nucleus and neutron-nucleus systems. As another important implication, exploring trade-offs of band-diagonal and full Hamiltonian matrices for different encodings is critical for the simplest two-cluster system of two nucleons; namely, this allows quantum simulations of nuclear structure to handle the complete form of the chiral nucleon-nucleon (NN) potentials (e.g., see [21, 24, 22]), which are in turn key to ab initio large-scale simulations of light, medium-mass, and even selected heavy nuclei.
[…]
@@
Our paper is intended to serve interdisciplinary research at the intersection of nuclear physics and quantum information science, and in some cases includes details that may be well known in one of the fields but are pedagogical for researchers of the other field: our aim here is to provide a complete framework for the problem at hand. Our paper is organized as follows. In Sec. II, we introduce the nuclear problem of solving the neutron-nucleus dynamics and its Hamiltonian. In Sec. III, we discuss different encoding methods of mapping the given Hamiltonian to a form that can be simulated on a quantum computer. The different encoding methods we discuss include the one-hot encoding, binary encoding, and the Gray encoding. In Sec. IV, we briefly explain the variational principle, for completeness of presentation. Since the variational principle depends on the choice of a trial state, called ansatz, we delineate the different ansatz choices for the different encodings. In Sec. V, we analyze various advantages between the different encodings considered, including the number of Pauli terms and the number of commuting sets, for a most general local potential and its band-diagonal approximation. As part of our trade-off analysis for the different encodings of Hamiltonians, we introduce the new DGC measurement scheme to group Pauli strings into commuting operator sets. For this scheme we provide the explicit diagonalizing unitary and an analysis of the number of commuting sets. In Sec. VI, we provide quantum simulations for different encodings and nuclear systems, including comparisons of the different commuting sets. We discuss the results and challenges for weakly bound states.
@@ II Problem description
$V(r)u(r)=\int dr^{\prime}\,r^{\prime 2}\,\tilde{V}(r,r^{\prime})\,u(r^{\prime}),$ (4)
where $u(r)$ is in units of fm-3/2 and $V(r)$ is in units of MeV. We note that for bound states and resonances for which only the elastic channel is open, the potentials are real, except on the poles, which can be numerically avoided through the principal value theorem as shown in Ref. [11]; while in this study, we do not need the imaginary part of the $\tilde{V}(r,r^{\prime})$ optical potential for the bound-state simulations at hand, the generalization to complex matrices is straightforward and feasible. The potential energy $V(r)$ enters into the two-body Schrödinger equation, as described next.
@@
where $n_{r}$ denotes a relative harmonic oscillator radial node number. The matrix elements of $T$ for every $\ell$ are given by
$\langle n_{r}^{\prime}\ell|T|n_{r}\ell\rangle\coloneqq\frac{\hbar\omega}{2}\Bigg[\bigg(2n_{r}+\ell+\frac{3}{2}\bigg)\delta_{n_{r}^{\prime},n_{r}}\\ \qquad-\sqrt{n_{r}\bigg(n_{r}+\ell+\frac{1}{2}\bigg)}\delta_{n_{r}^{\prime},n_{r}-1}\\ \qquad-\sqrt{\bigg(n_{r}+1\bigg)\bigg(n_{r}+\ell+\frac{3}{2}\bigg)}\delta_{n_{r}^{\prime},n_{r}+1}\Bigg],$ (9)
@@
$\langle n_{r}^{\prime}\ell|r^{2}|n_{r}\ell\rangle\coloneqq b_{s}^{2}\Bigg[\bigg(2n_{r}+\ell+\frac{3}{2}\bigg)\delta_{n_{r}^{\prime},n_{r}}\\ \qquad+\sqrt{n_{r}\bigg(n_{r}+\ell+\frac{1}{2}\bigg)}\delta_{n_{r}^{\prime},n_{r}-1}\\ \qquad+\sqrt{\bigg(n_{r}+1\bigg)\bigg(n_{r}+\ell+\frac{3}{2}\bigg)}\delta_{n_{r}^{\prime},n_{r}+1}\Bigg],$ (12)
@@
The quantum computational simulations of nuclear systems of Ref. [18, 16] have used a contact potential for $\ell=0$, or $\left\langle n_{r}^{\prime}\right|V\left|n_{r}\right\rangle=V_{0}\delta_{n_{r},n_{r}^{\prime}}\delta_{n_{r},0}$ (Fig. 1a). In this work, the Hamiltonian is generalized for any band-diagonal to full matrix and can now accommodate a general central potential, such as exponential Gaussian-like potentials using Eq. (13) (e.g., [15]), along with ab initio inter-cluster potentials (as those derived in Ref. [11]) and the central part of chiral NN potentials using Eq. (7). In general, chiral NN potentials, such as in Refs. [21, 24, 22], used in ab initio nuclear calculations require, in addition, spin and isospin degrees of freedom $\alpha=n_{r}(\ell\frac{1}{2})jmt_{z}$, as described above. Including the additional spin-isospin quantum numbers leads to a larger set of basis states, but nonetheless, in the present framework this is straightforward by generalizing Eq. (8) to
@@
where $\langle\alpha^{\prime}|V|\alpha\rangle$ are known matrix elements for any chiral NN potential and the enumerating index $\alpha$ replaces $n_{r}$ in the mappings to qubits discussed next. Importantly, physically relevant potentials with a comparatively soft core or widely used potentials renormalized using, e.g., the Similarity Renormalization Group (SRG) technique [8], are band-diagonal as a result of the decoupling of low- and high-momentum configurations. Hence, the advantages found in this study as a function of the bandwidth $2K+1$ directly generalize to the complete form of the chiral NN potentials and their routinely used band-diagonal structure.
@@ III Mapping onto a quantum computer
Mappings from operators in the Fock space to Pauli strings are called encodings. Multiple encodings can be found in existing literature, for example, the one-hot encoding (OH), the Bravyi-Kitaev encoding, the Verstraete-Cirac encoding, the binary encoding, and the Gray encoding (see, e.g., [43, 16, 53]. In addition, there has been substantial work on using qudits (higher dimensional systems) rather than qubits, which has been shown to provide some advantage in specific scenarios [28, 55, 7]. In this work, we focus on three qubit encodings - OH, binary, and Gray - and analyze various properties and trade-offs.
To provide an illustrative example, for each encoding above, we show the explicit Pauli terms for the $N=4,K=2$ Hamiltonian with the exponential potential in Eq. (13) with parameters
$V_{0}=-2.79\,\textrm{MeV},\,c=0.05,\,\hbar\omega=15.95\,\textrm{MeV},$ (17)
for the case of n+16C with reduced mass $\mu=\frac{16}{17}m_{\textrm{N}}$ (here we use $m_{\textrm{N}}=938.272029$ MeV for both protons and neutrons). The Hamiltonians for the different encodings can be found in their respective sections below.
@@ III.0.1 One-hot encoding
$|0\rangle$ $\to$ $|1000\rangle$,
$|1\rangle$ $\to$ $|0100\rangle$,
$|2\rangle$ $\to$ $|0010\rangle$,
$|3\rangle$ $\to$ $|0001\rangle$.
Next, we define the number operators and ladder operators. The number operator $|m\rangle\!\langle m|$ maps $|m\rangle$ to itself and maps all other basis states to 0:
@@
In this encoded basis, this action involves flipping the $(m\pm i)$ qubit to 0 and the $m$ qubit to 1:
@@
The upper limit of the $k$ summation is $\max(K,1)$ due to the fact that even if $K=0$, the first off-diagonal term in the Hamiltonian is non-zero as a result of the kinetic energy having two off-diagonal terms. Note that, to preserve the size of the matrix to be $N$, the summation over $k$ terminates if m+k>N1m+k>N-1. Thus, in the encoded one-hot basis, if $K>1$, the resulting matrix is $(2K+1)$-diagonal.